1. Rikell company produces helmets using labor (L) and capital (K). Its production function is given...

1.

Rikell company produces helmets using labor (L) and capital (K). Its production function is given by the following expression:

Q = 98 L + 5 K

where Q is the output of helmets.

The prices of labor (PL), capital (PK), helmet (P) and the cost (C) are the following:

PL=70, PK=62, P=48 and C=2,856

What is the slope of Rikell's Isoquant?

2.

Rikell company produces helmets using labor (L) and capital (K). Its production function is given by the following expression:
Q = 37 L + 57 K

where Q is the output of helmets.

The prices of labor (PL), capital (PK), helmet (P) and the cost (C) are the following:
PL=10,   PK=25 P=90   C =4,767
What is the slope of the Isocost?
Answer #1

1. Slope of the isoquant = -MRTS
MRTS = Marginal product of labor/Marginal product of capital =

So, slope of isoquant = -19.6

2. Slope of the isocost = -PL/PK = -10/25 = -0.4
So, slope of the isocost is -0.4